The Euro return (capital gains and dividend yield) on XXX share was 15% between 1st Jan 2016 and 31st Dec 2016. The dollar return on XXX share was 20.58% over the same period. The dollar/Euro exchange rate was 1.3 on 1st Jan 2016. The annual dividend yield for XXX share was 7% on 2nd Jan 2016.

a) On 31st Dec 2016, what exchange rate would you expect? What would have most likely happened to the Dollar over the one-year period (1st Jan 2016 – 31st Dec 2016)? Explain your calculation )

(b) Ignore part (a). If the share price for XXX was Euro100 on 1st Jan 2016, what would the XXX share price most likely to be on 31st Dec 2016?
Answer #1

(a) Eur Return = 15 % , USD Return = 20.58%, Dividend Yield = 7%

Capital Gains Yield in EUR = 15 - 7 = 8%

Capital Gains Yield in USD = 20.58 - 7 = 13.58 %

Initial Exchange Rate = $ 1.3 / EUR

Let the final Exchange Rate be $ K / EUR

Now let the initial investment be EUR 1

Therefore, Initial $ Investment = 1.3 x 1 = $ 1.3

Final Value of EUR Investment = 1.08 x 1 = EUR 1.08

Final Value of $ Investment = 1.1358 x 1.3 = $ 1.47654

$ Return = [(1.08 x K) - 1.3] / 1.3 = 0.1358

1.08 x K = 1.47654

K = 1.47654 / 1.08 = $ 1.367 ~ $ 1.37/ EUR

The $ Value depreciates against the EUR value over the 1-Year period as the $ Interest Rate is greater than the EUR Interest Rate.

(b) If the Expected Capital Gains Yield in EUR is 8% and the Initial price is EUR 100, then the final price is 100 x 1.08 = 108 EUR